# Visual Whole-Body Control for Legged Loco-Manipulation

### v2

@@ I Introduction
While this is appealing, it is a very challenging control problem to coordinate all the joints simultaneously with large degrees of freedom (19 DoF). The robot will need to exploit the contact with its surroundings and the objects and maintain stability and robustness to external disturbances all at the same time. An even larger challenge comes from achieving all these in diverse environments autonomously, given only the observation from an egocentric camera. Recent learning-based approaches have shown promising results on legged robots avoiding obstacles, climbing stairs, and jumping over stages robustly with visual inputs [13, 1, 78, 15]. However, all these efforts still focus only on locomotion without manipulation, which requires more precise control. The extra complexity and more challenging tasks make direct end-to-end learning infeasible.
@@
Our hardware platform is built on a Unitree B1 quadruped robot equipped with a Unitree Z1 robotic arm. We design a server-client multi-processing system that consists of four different modules working under different frequencies. We conducted the pickup task with 14 different objects, including regular shapes, daily uses, and irregular objects, under three height configurations, i.e., on the ground, on a box, and a table, achieving a high success rate in all scenarios.
@@ II Related Work
Legged Locomotion. For decades, legged locomotion has been an essential topic in the field of robotics and has gained much success. Classical methods solve the legged locomotion task through model-based control to define controllers in advance. For example, Miura and Shimoyama [49], Sreenath et al. [58] utilized traditional methods to enable the walking of biped robots. MIT cheetah 3 [7], ANYmal-a [31] by ETH, and Atlas [19] by Boston Dynamics are all perfect examples of robots designed with a robust walking controller. Based on these developed robots, researchers have been able to develop advanced controllers for even more agile locomotion tasks, such as jumping or walking over varying obstacles [52, 50, 25, 5, 51, 3, 38]. To make a robust controller that can adapt to varying environments, researchers integrate some sensor data such as elevation maps [18, 40, 35, 6, 17] and state estimators [8, 63, 10, 70, 62].
@@
Although classical methods have greatly succeeded in legged locomotion, generalizability is always a problem for pre-designed controllers, and acquiring these controllers requires careful and tedious engineering. With the fast development of deep neural networks, learning-based end-to-end methods have become a simple way to gain robust controllers with strong generalization abilities. Learning-based methods often train a robust controller in physics simulators such as PyBullet [14] and Isaac Gym [46] and then perform sim-to-real transfer to a real robot. For example, Rudin et al. [56] uses Isaac Gym with GPU to enable a quadruped robot to walk in only several minutes, Kumar et al. [36, 37] proposed the RMA algorithm to make the sim-to-real transfer process easier, and more robust, Lee et al. [41], Miki et al. [48], Agarwal et al. [1], Yang et al. [69], Yu et al. [76], Cheng et al. [13], Zhuang et al. [78], Fu et al. [20], Margolis and Agrawal [47], Li et al. [44] enable the legged robots to walk or run on varying terrains, and Fuchioka et al. [23], Escontrela et al. [16], Peng et al. [53], Yang et al. [67], Wang et al. [61], Li et al. [43], Kumar et al. [39] combine RL with imitation learning to produce naturalistic and robust controllers for legged robots. Kim et al. [33] map human motions to a quadruped robot to complete a wide range of locomotion and manipulation tasks.
@@
Mobile Manipulation. Previously mentioned works mainly focus only on the mobility part; adding manipulation to the mobile base is yet to be well-solved. Many recent efforts have focused on building teleoperation systems for mobile manipulation [60, 64, 21, 22]. For example, Fu et al. [21] proposed an effective method to train a whole-body controller end-to-end through DRL. However, the controller requires human involvement to conduct different tasks instead of operating autonomously. Autonomous Mobile manipulation with wheels has made significant progress over the past few years [27, 9, 2, 57, 68]. Lew et al. [42] built an efficient framework for table cleaning with RL and trajectory optimization. [24] and Srivastava et al. [59] integrate task-specific planner for various mobile manipulation tasks. Gu et al. [26] and Xia et al. [65] developed modularized mobile manipulators with RL, Yokoyama et al. [74] introduced a high-level controller for coordinating different low-level skills. Yang et al. [68] proposed simultaneously controlling the robot arm and base for harmonic mobile manipulation using RL training with large-scale environments. While encouraging, most mobile manipulators with wheels cannot operate with challenging terrains in the wild. Zhang et al. [77], Yokoyama et al. [75], Arcari et al. [4] all performed an autonomous mobile manipulation system on quadruped robots while they utilized the default low-level controller and only focused on the manipulation part without any whole-body behaviors to adapt to different heights of objects.
@@
The system we design highly differs from the aforementioned works, as our system is a fully autonomous legged mobile manipulation system. The whole robot is controlled and planned by trained policies and can adapt to objects with varying heights through its whole-body behavior. In Fig. 2 we show why VBC is highly advantaged of the non-whole-body methods [77, 75, 4], for the body control behavior allowing much more flexible poses (e.g., bending) to reach objects in random heights, like the one on the ground.
@@ III-B Low-Level Policy for Whole-Body Goal-Reaching
The low-level control policy $\pi^{\text{low}}$ as described in blue part in Fig. 3 is for whole-body locomotion control and can provide a robust controller that allows the robot to track any given end-effector poses and velocities with its full DOF robustly across various terrains. In this work, we turn to the help of RL and utilize ROA (Regularized Online Adaptation) [21, 12, 32] to realize such a universal behavior that can become the foundation of any high-level tasks.
@@ C.1. Privileged Teacher Policy
Figure 6: The server-client multi-processing system design. Modules work under different frequencies.
@@
Domain randomization. When training our high-level policy, we randomized the friction between the robot and the terrain, the mass of the robot, and the center of mass of the robot. Although our low-level control policy can handle these gaps in the real world, its performances differ under different environmental conditions. These randomizations make the high-level policy adjust to different low-level performances.
@@ C.2. Visuomotor Student Policy
Imitation learning for student distillation. To deploy our policy in the real world, we distill our state-based high-level planning policy into a vision-based policy using DAgger. We warm up by teacher sampling the initial transition data. This is like behavior cloning (BC) to find a good initialization of the visuomotor student policy to improve the training efficiency. Then we sample transitions using the student policy and request correction from the teacher. More details can be found in Appendix I.
Policy network architecture. Our vision-based high-level policy network comprises a CNN, a GRU, and a two-layer MLP that outputs the high-level actions. We concatenate the images and other states as observations. The images are passed through the two-layer CNN network with kernel size being $5$ and $3$, respectively, and then encoded as a $64$-dim latent vector. Then, we concatenate this latent vector with other states as a flattened vector. This vector is then passed to a one-layer GRU with $128$ hidden size, followed by a two-layer MLP with hidden size $64$. We use ELU as the activation.
@@ III-D Real World Deployment
System design. During deployment, the low-level policy that controls the quadruped works at a frequency of 50Hz; in the meantime, the high-level policy runs at 10Hz, which is five times slower than the low-level policy; besides, the arm is required to work at a frequency of more than 500Hz. Since these policies work on different computers, along with the quadruped and the arm should work under different frequencies, we design a client-server multi-processing system for this work, as shown in Fig. 6. We set up three processes at the low-level policy side: a quadruped robot control process, an arm control process, and a goal receiver process behaving as a client. The high-level policy combines the vision inputs from the depth cameras and the information from the low-level request as the observation works as a server. Every 0.1 seconds, the goal receiver sends a request with the necessary information to the server; the high-level policy makes the latest inference and sends the high-level action back to the robot; at the same time, the quadruped and the arm control processes gather the latest high-level commands to execute, while keep updating the necessary information for the goal receiver process.
@@
Vision masks. As we need to segment objects from environment backgrounds, we develop TrackingSAM, a tool that combines the tracking model AOT (Associating Objects with Transformers) [72] and SAM (Segment Anything) [34]. This allows us to annotate the object on the image from both cameras or an arbitrary image before starting the task and track that object to provide its real-time segmented masks. The tool is integrated into the high-level server, and a demo example can be found in Fig. 7. During real-world experiments, we annotated the objects from the two RGBD cameras mounted on the gripper and the robot head at the beginning of each reset trial. TrackingSAM will keep tracking the segmented object and providing the mask. The depth images obtained from the RGBD cameras are pre-processed by hole-filling and clipping.
@@ IV Experiments
Figure 10: Success rates of different student policies at different heights over three seeds, tested in the simulator.
@@ IV-A Simulation Results
Evaluation principles and baselines. In the simulation, we test the picking-up success rate on each distinct object over $300$ episodes and collect the results over every category.. When the environment is reset, it’s considered a success if an object is picked up and a failure when not successful in 150 high-level steps. To be specific, we compare the final visuomotor student policy combined with the whole-body low-level policy trained by VBC to two visuomotor baseline methods over $34$ kinds of different objects: 1) VBC w.o. shape feature: a visuomotor policy that is trained the same as VBC, but does not access the pre-trained object shape features during the teacher training stage. This aims to show the necessity of the object shape features during teacher training. 2) Floating base: a floating base policy with a perfect low-level navigation ability but without whole-body behavior; in other words, the robot can always follow the given velocity commands at a fixed body height (0.55m in our case). This is to show the indispensability of the whole-body behavior. During test time, all object heights are randomly set from $0.0$m to $0.6$m for each trial. 3) Non-hierarchical: a unified policy trained with low-level policy and visuomotor high-level policy jointly in an end-to-end style. This policy takes the observation of our low-level and high-level policies together. It outputs all the target positions of 12 robot joint angles and the target pose of the gripper. We failed in training such a policy, demonstrating that the training pipeline of VBC is inevitable.
TABLE I: Success rates of VBC compared to baseline method on seven categories of objects, tested in the simulator. Results are evaluated over three seeds.
@@
VBC (Ours) $\mathbf{70.51\%}$ $\mathbf{51.40\%}$ $76.53\%$ $\mathbf{69.79\%}$ $\mathbf{81.22\%}$ $\mathbf{74.16\%}$ $\mathbf{56.70\%}$
VBC w.o. Shape Feature $68.44\%$ $33.88\%$ $\mathbf{77.72\%}$ $54.18\%$ $65.42\%$ $17.95\%$ $9.04\%$
Floating Base $49.75\%$ $32.71\%$ $56.38\%$ $45.37\%$ $25.36\%$ $27.18\%$ $12.62\%$
Non-Hierarchical $0.0\%$ $0.0\%$ $0.0\%$ $0.0\%$ $0.0\%$ $0.0\%$ $0.0\%$
@@
Figure 11: Average success times w.r.t. the trial times of picking up 14 objects at different heights, tested in the real world. A failure is recorded if the robot attempts more than 5 trails in one reset or if the object falls from a higher surface.
Picking up different objects. As is shown in Tab. I, VBC achieves the best performance across almost all categories of objects, especially on objects with the most complicated shapes, such as bowl and drill. One may notice that VBC exhibits small differences from VBC w.o. pre-trained shape features in objects like ball and square box. The reason is that, without pre-trained features, the policy targets the object’s center for grasping, which is effective enough on simpler objects yet inapplicable for objects with more complicated shapes. Therefore, we conclude that the shape feature does help a lot in picking up objects in general. In addition, the performance of all policies behaves the worst on long boxes, as the long box has several inaccessible poses under the current gripper when we randomize the initial pose during experiments.
@@
Picking up over various heights. We count the success rate of VBC and the floating base baseline on seven different heights and illustrate the results in Fig. 10. From the significant improvement of VBC, compared to the floating base policy whose heights are fixed, we know why such a hierarchical method of VBC should be applied to achieve flexible whole-body behaviors in achieving high-level tasks.

### v3

@@ I Introduction
While this is appealing, it is a very challenging control problem to coordinate all the joints simultaneously with large degrees of freedom (19 DoF). The robot will need to exploit the contact with its surroundings and the objects and maintain stability and robustness to external disturbances all at the same time. Besides, acquiring a mobile robot platform to do precise manipulation jobs requires a steady robot body, which is much harder for legged robots compared to wheeled robots. An even larger challenge comes from achieving all these in diverse environments autonomously, given only the observation from an egocentric camera. Recent learning-based approaches have shown promising results on legged robots avoiding obstacles, climbing stairs, and jumping over stages robustly with visual inputs [13, 1, 82, 15]. However, all these efforts still focus only on locomotion without manipulation, which requires more precise control. The extra complexity and more challenging tasks make direct end-to-end learning infeasible.
@@
Our hardware platform is built on a Unitree B1 quadruped robot equipped with a Unitree Z1 robotic arm. We design a server-client multi-processing system that consists of four different modules working under different frequencies. We conducted the pickup task with 14 different objects, including regular shapes, daily uses, and irregular objects, under three height configurations, i.e., on the ground, on a box, and on a table, achieving a high success rate in all scenarios.
We summarize the contribution of this paper as:
  • We develop an autonomous vision-based loco-manipulation system that utilizes model-free RL and sim-to-real techniques to learn whole-body control policies for legged robots.
  • Our proposed method emerges retrying behavior and generalizes well to complicated terrains, various objects, and heights.
@@ II Related Work
Legged Locomotion. For decades, legged locomotion has been an essential topic in the field of robotics and has gained much success. Classical methods solve the legged locomotion task through model-based control to define controllers in advance. For example, Miura and Shimoyama [50], Sreenath et al. [61] utilized traditional methods to enable the walking of biped robots. MIT cheetah 3 [7], ANYmal-a [31] by ETH, and Atlas [19] by Boston Dynamics are all perfect examples of robots designed with a robust walking controller. Based on these developed robots, researchers have been able to develop advanced controllers for even more agile locomotion tasks, such as jumping or walking over varying obstacles [53, 51, 26, 52, 3, 38]. To make a robust controller that can adapt to varying environments, researchers integrate some sensor data such as elevation maps [18, 40, 35, 6, 17] and state estimators [8, 66, 10, 74, 65]. Although classical methods have greatly succeeded in legged locomotion, generalizability is always a problem for pre-designed controllers, and acquiring these controllers requires careful and tedious engineering. With the fast development of deep neural networks, learning-based end-to-end methods have become a simple way to gain robust controllers with strong generalization abilities. Learning-based methods often train a robust controller in physics simulators such as PyBullet [14] and Isaac Gym [47] and then perform sim-to-real transfer to a real robot. For example, Rudin et al. [58] uses Isaac Gym with GPU to enable a quadruped robot to walk in only several minutes, Kumar et al. [36, 37] proposed the RMA algorithm to make the sim-to-real transfer process easier, and more robust, Lee et al. [41], Miki et al. [49], Agarwal et al. [1], Yang et al. [73], Yu et al. [80], Cheng et al. [13], Zhuang et al. [82], Fu et al. [21], Margolis and Agrawal [48], Li et al. [44] enable the legged robots to walk or run on varying terrains, and Fuchioka et al. [24], Escontrela et al. [16], Peng et al. [54], Yang et al. [71], Wang et al. [64], Li et al. [43], Kumar et al. [39] combine RL with imitation learning to produce naturalistic and robust controllers for legged robots. Kim et al. [33] map human motions to a quadruped robot to complete a wide range of locomotion and manipulation tasks.
@@
Mobile Manipulation. Previously mentioned works mainly focus only on the mobility part; adding manipulation to the mobile base is yet to be well-solved. Autonomous mobile manipulation on wheeled robot platforms has made significant progress over the past few years. Wong et al. [68] collected teleoperated mobile manipulation data and learned imitation-learning-based policies in simulation. Lew et al. [42] built an efficient framework for table cleaning with RL and trajectory optimization. Garrett et al. [25] and Srivastava et al. [62] integrate task-specific planner for various mobile manipulation tasks. Gu et al. [27] and Xia et al. [69] developed modularized mobile manipulators with RL, Yokoyama et al. [78] introduced a high-level controller for coordinating different low-level skills. Sun et al. [63] designed a modularized policy with components for manipulation and navigation, trained by RL to learn navigation and grasping indoors. Yang et al. [72] proposed simultaneously controlling the robot arm and base for harmonic mobile manipulation using RL training with large-scale environments. Ahn et al. [2] pre-trained a set of single mobile or manipulation tasks and relied on large language models to decompose and call the skills to achieve long-horizon tasks. Brohan et al. [9] train a robotics transformer to perform real-world tasks given human instructions, trained on large-scale and real-world demonstration. Shafiullah et al. [59] learned to solve household tasks from 5-minute demonstrations based on pre-trained home representations. While encouraging, most of these mobile manipulators with wheels cannot operate with challenging terrains in the wild.
Legged Loco-Manipulation A set of works has exploited legged loco-manipulation to take advantage of legged robots. Bellicoso et al. [5] designs an online motion planning framework, together with a whole-body controller based on a hierarchical optimization algorithm to achieve loco-manipulation, Fu et al. [22] proposed an effective method to train a whole-body controller end-to-end through DRL. Ma et al. [46] combined model-based manipulator control and learning-based locomotion, but only showed the response to external push instead of achieving manipulation tasks. Notably, these works all require human involvement and teleoperation to execute and achieve tasks, rather than operating autonomously. Zhang et al. [81], Yokoyama et al. [79], Arcari et al. [4] all performed an autonomous mobile manipulation system on quadruped robots while they utilized the default model-based controller provided through the robot APIs (only control robot locomotion in a static height) and only focused on the manipulation part without any whole-body behaviors to adapt to different heights of objects. Ferrolho et al. [20] proposed a trajectory optimization framework based on a robustness metric for solving complex loco-manipulation tasks, but the task is assigned and the target positions are calculated via a motion capture system, thus only works limitedly. Sleiman et al. [60] realized a bilevel search strategy for motion planning, by incorporating domain-specific rules, combing trajectory optimization, informed graph search, and sampling-based planning, enabling robots to perform complex tasks. They require the user to provide a dense description of the scene to initialize the solver. Zimmermann et al. [83] also utilized trajectory optimization to enable a legged robot to walk and pick up objects. However, they only show limited surroundings and simple objects. Besides these works with arm-based configuration, Wolfslag et al. [67] use the legs for pushing and box lifting tasks by introducing additional support legs named prongs to enhance the robustness and manipulation capabilities of quadrupedal robots. Jeon et al. [32] adopted a similar hierarchical framework to us, but does not involve any visual inputs, and only learned a quadruped for pushing large objects without an additional arm.
@@
The system we design highly differs from the aforementioned works, as our work is a fully autonomous and vision-based mobile manipulation system on quadruped robots, with limited human intervention. The robot is controlled and planned by trained policies and can adapt to objects with varying heights through its whole-body behavior, and emergent retrying behaviors when the task fails. Due to our design of visual inputs, our system can work both indoors and outdoors with no external limits, as illustrated in Fig. 1, compared with almost all of the above-mentioned works that are only tested in limited lab scenarios. In Fig. 2 we show why VBC is highly advantaged of the non-whole-body methods [81, 79, 4], for the body control behavior allowing much more flexible poses (e.g., bending) to reach objects in random heights, like the one on the ground.
@@ III-B Low-Level Policy for Whole-Body Goal-Reaching
The low-level control policy that takes over the whole-body locomotion control $\pi^{\text{low}}$ as described in the blue part in Fig. 3, is trained to track any given end-effector poses and body velocities across various terrains. In this work, we turn to the help of RL and utilize ROA (Regularized Online Adaptation) [22, 12, 32] to realize such a universal and robust behavior that can become the foundation of any high-level tasks.
@@ C.1. Privileged Teacher Policy
Figure 6: The server-client multi-processing system design. Modules work asynchronously under different frequencies. The high-level planning policy (left) and the low-level control modules (right, controlling arm and legs) are deployed on different devices. To be more specific, the goal receiver requests the high-level policy to take the visual input and receive the new commands; the low-level policy sends joint position commands to the quadruped runs at 50Hz, given the input of high-level commands; the IK-based arm controller runs and send the joint position command to the arm constantly at 800Hz, given the target end-effector pose. Therefore, the low-level policy and the arm controller keep utilizing the old commands to control the robot until the high-level policy updates the commands (i.e., target robot velocities and end-effector pose).
@@
Domain randomization. When training our high-level policy, we randomized the friction between the robot and the terrain, the mass of the robot, and the center of mass of the robot. Although our low-level control policy can handle these gaps in the real world, its performances differ under different environmental conditions. These randomizations help the high-level policy adjust better to different low-level performances. Moreover, due to the uncertainty of the high-low signal transmission in the real world, we also randomize the high-level frequency by involving different low-level policy calls during one high-level call. We also randomize the Kp and Kd parameters for the arm controller as we use the built-in position control mode for the arm in the simulation. Regarding the pick-up task, we randomize table heights, the initial position and pose of the robot, along with the position and pose of objects.
@@ C.2. Visuomotor Student Policy
Imitation learning for student distillation. To deploy our policy in the real world, we distill our state-based high-level planning policy into a vision-based policy using DAgger. We warm up by teacher sampling the initial transition data. This is like behavior cloning (BC) to find a good initialization of the visuomotor student policy to improve the training efficiency. Then we sample transitions using the student policy and request correction from the teacher. We adopt augmentations like RandomErasing, GaussianBlur, GaussianNoise, and RandomRotation for the depth images. More details can be found in Appendix II-B.
Policy network architecture. Our vision-based high-level policy network comprises a CNN, a GRU, and a two-layer MLP that outputs the high-level actions. We concatenate the images and other states as observations. The images are passed through the two-layer CNN network with kernel size being $5$ and $3$, respectively, and then encoded as a $64$-dim latent vector. Then, we concatenate this latent vector with other states as a flattened vector. This vector is then passed to a one-layer GRU with hidden size 128, followed by a two-layer MLP with hidden size 64. We use ELU as the activation.
@@ III-D Real World Deployment
System design. During deployment, the low-level policy that controls the quadruped works at a frequency of 50Hz; in the meantime, the high-level policy runs at 8Hz, which is five times slower than the low-level policy; besides, the arm is required to work at a frequency of more than 800Hz. Since these policies work on different computers, along with the quadruped and the arm should work under different frequencies, we design a client-server multi-processing system for this work, as shown in Fig. 6. We set up three processes at the low-level policy side: a quadruped robot control process, an arm control process, and a goal receiver process behaving as a client. The high-level policy combines the vision inputs from the depth cameras and the information from the low-level request as the observation works as a server. Every 0.1 seconds, the goal receiver sends a request with the necessary information to the server; the high-level policy makes the latest inference and sends the high-level action back to the robot; at the same time, the quadruped and the arm control processes gather the latest high-level commands to execute, while keep updating the necessary information for the goal receiver process.
@@
Vision masks. Note that our system works with only limited human intervention, i.e., we need to annotate and segment objects from environment backgrounds. To this end, we develop TrackingSAM [56], a tool that combines the tracking model AOT (Associating Objects with Transformers) [76] and SAM (Segment Anything) [34]. This allows us to annotate the object on the image from both cameras or an arbitrary image before starting the task and track that object to provide its real-time segmented masks. The tool is integrated into the high-level server, and a demo example can be found in Fig. 7. During real-world experiments, we annotated the objects from the two RGBD cameras mounted on the gripper and the robot head at the beginning of each reset trial. TrackingSAM will keep tracking the segmented object and providing the mask. The depth images obtained from the RGBD cameras are pre-processed by hole-filling and clipping. In real-world experiments, as long as the object is visible to one camera (either the head or the wrist camera), the annotation mask can be propagated to both views to initialize an automated pickup.
@@ IV Experiments
Figure 10: Success rates of different student policies at different heights, tested in the simulator. Each distinct object for more than 100 episodes. The robot starts from a 1.0-meter distance from the table. Every episode ends when the robot successfully picks up the object, the object falls from the table, or the robot fails until 150 steps ($\sim$ 18 seconds). The dots represent the mean performance of all objects, the upper and the lower bound represent the maximum and the minimum success rate overall 7 categories, where we also note their category name.
@@ IV-A Simulation Results
TABLE I: Success rates of VBC compared to baseline method on seven categories of objects, tested in the simulator. Each distinct object for more than 300 episodes. The robot starts from a 1.0-meter distance from the table. Every episode ends when the robot successfully picks up the object, the object falls from the table, or the robot fails until 150 steps. With the 3D shape feature, the algorithm is better at handling irregular objects. Detailed descriptions of each method can be referred to Section IV-A.
@@
Privleged Teacher Floating Base $63.25\%$ $20.51\%$ $72.66\%$ $38.07\%$ $69.55\%$ $71.88\%$ $4.17\%$
VBC w.o. Shape Feature $\mathbf{96.18}\%$ $74.26\%$ $\mathbf{96.80}\%$ $\mathbf{82.46}\%$ $84.38\%$ $55.84\%$ $73.77\%$
VBC (Ours) $89.08\%$ $\mathbf{92.83\%}$ $81.87\%$ $76.34\%$ $\mathbf{89.12\%}$ $\mathbf{78.14\%}$ $\mathbf{84.00\%}$
Visuomotor Student Non-Hierarchical $0.0\%$ $0.0\%$ $0.0\%$ $0.0\%$ $0.0\%$ $0.0\%$ $0.0\%$
Floating Base $42.45\%$ $0.0\%$ $37.36\%$ $8.33\%$ $41.41\%$ $43.57\%$ $3.70\%$
VBC (Ours) $\mathbf{55.40}\%$ $\mathbf{28.57}\%$ $\mathbf{80.00}\%$ $\mathbf{56.57}\%$ $\mathbf{68.01}\%$ $\mathbf{58.96\%}$ $\mathbf{53.33\%}$
@@
Figure 11: Average success times w.r.t. the trial times of picking up 14 objects at different heights, tested in the real world. We allow 5 continuing trial times (including 4 continuous retries, x-axis) during one experiment, i.e., one reset of picking up one specific object. A failure is recorded if the robot attempts more than 5 trails in one reset or if the object falls from a higher surface. For each object and each height, such experiments are conducted 5 times, and therefore we calculate the average success time (y-axis) across these trials, representing the mean number of success times across five trials. It is obvious that the emergent retrying behavior improves the performance.
Evaluation principles and baselines. In the simulation, when the environment is reset and a new trail begins, both the robot and the objects are initialized with random positions and poses, and it’s considered a success if an object is picked up and a failure when the object falls down the table or is not successful in 150 high-level steps. During experiments, we compare both the privileged teacher policies of VBC against baselines, and their visuomotor student policies over $34$ kinds of different objects: 1) VBC w.o. shape feature: a visuomotor policy that is trained the same as VBC, but does not access the pre-trained object shape features during the teacher training stage. This aims to show the necessity of the object shape features during teacher training. We only compare the teacher’s performance as the only difference comes from the teacher’s side. 2) Floating base: a floating base policy with a perfect low-level navigation ability but without whole-body behavior; in other words, the robot can always follow the given velocity commands and body height commands ranging from 0.4 to 0.55 meters (we set 0.4 meters as the lower bound as the default controller in real-world only allows 0.47 meters in minimum). This is to show the indispensability of the whole-body behavior. We compare both the teacher and its corresponding student policy using the same policy network structure. 3) Non-hierarchical: a unified policy trained with low-level policy and visuomotor high-level policy jointly in an end-to-end style. This policy takes the observation of our low-level and high-level policies together. It outputs all the target positions of 12 robot joint angles and the target pose of the gripper. We failed in training such a policy, demonstrating that the training pipeline of VBC is inevitable.
Figure 12: Visualization of robot behavior and corresponding commands during the test time. Particularly, we plot two kinds of high-level commands given by the high-level policy, as described in Section III-C, i.e., the velocity commands (including the quadruped linear velocity and yaw velocity); along with the gripper command (indicating whether the gripper should be close or open).
Picking up different objects. We test the picking-up success rate on each distinct object for more than $300$ episodes and collect the results over every category. During test time, all object heights are randomly set from $0.0$m to $0.6$m for each trial. As is shown in Tab. I, VBC achieves the best performance on 4/7 categories of objects. It is also interesting that with 3D features, VBC becomes worse at regular objects (compared to VBC w.o. Shape Feature), but is much better on irregular objects with complicated shapes, i.e., long box, cup, bowl and drill. One possible reason is that, without pre-trained features, the policy targets the object’s center for grasping, which is effective enough on simpler objects yet inapplicable for objects with more complicated shapes. One may notice that the student policy is much worse than the teacher policy, this is because when the robot touches but fails to grasp on the first try, the object will be randomly laid so that its pose will be much harder to grasp within only 150 steps using only visual inputs. In addition, the performance of the student policy of VBC behaves the worst on long boxes. This is because without shape features, it is much harder to determine the pose of the object, and some of the initial poses of long boxes are extremely hard to grasp for the gripper we used at some height (also as shown in Fig. 10).
@@
Picking up over various heights. We test the mean success rate on each object category on a fixed height, where we also list the object with the maximum/minimum success rate on each height. We compare VBC and the floating base baseline on 7 different heights, and illustrate the results in Fig. 10. From the significant improvement of VBC on almost every height, compared to the floating base policy, we highlight the advantage of VBC that is able to achieve flexible whole-body behaviors in achieving high-level tasks.